# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@ II.1.1 Eigenstate solver
In this study, we ensured that our calculations produced wavefunctions $|\psi^{(k)}\rangle$ that are eigenfunctions of Hamiltonian symmetries and in particular singlet states, i.e., eigenfunctions of total spin with eigenvalue 0, $\hat{S}^{2}|\psi^{(k)}\rangle=0$. Conservation of symmetries is notoriously important because molecular eigenstates are joint eigenfunctions of $\hat{H}$ and its symmetries, including $\hat{S}_{z}$, $\hat{N}$, molecular point-group symmetries (if any), and $\hat{S}^{2}$. Conservation of $\hat{S}_{z}$ and $\hat{N}$ (and molecular point-group symmetries that are isomorphic to $\mathbb{Z}_{2}^{\times n}$) can be achieved in a relatively easy way, since the eigenfunctions of these operators are Slater determinants (SD) $|{\bf{x}}\rangle$, e.g.
$\hat{N}|{\bf{x}}\rangle=(N_{\alpha{\bf{x}}}+N_{\beta{\bf{x}}})|{\bf{x}}\rangle\;,\;\hat{S}_{z}|{\bf{x}}\rangle=(N_{\alpha{\bf{x}}}-N_{\beta{\bf{x}}})|{\bf{x}}\rangle\;.$ (S6)
@@
Therefore, in this work, we ensure that the configurations used to span the SCI subspace all have the desired eigenvalue of $\hat{S}_{z}$ and $\hat{N}$ as part of the configuration recovery procedure in the presence of noise. Conservation of $\hat{S}^{2}$ is more difficult to achieve in CI methods since the eigenfunctions of $\hat{S}^{2}$ are not SDs. The proper way of ensuring conservation of $\hat{S}^{2}$ is expanding $|\psi^{(k)}\rangle$ on a set of configuration state functions (CSFs), i.e. spin-symmetry-adapted linear combinations of SDs. The use of CSFs was common in early CI codes since, in addition to the obvious benefits of reducing the memory footprint of the CI vector, automatic conservation of total spin enhanced the stability of the CI iterations. Modern CI codes tend to employ SDs as opposed to CSFs, because the formation of the $\sigma$ vector (the most memory- and rate-limiting step of CI algorithms) is considerably more efficient and easily parallelized in SD-based algorithms, and in part because system memory and disk are more plentiful than in previous machines. However, in SD-based CI codes, the conservation of total spin is no longer guaranteed. In the case of our HPC quantum estimator, sampling from a quantum computer may return sets of SDs that do not allow constructing eigenfunctions of total spin. For example, in a (2e,2o) system, one may sample the configuration $|1001\rangle$ (having a single spin-down excitation over the RHF state $|0101\rangle$) which is a linear combination of the open-shell singlet and triplet states, respectively $\left(|1001\rangle\pm|0110\rangle\right)/\sqrt{2}$. If the configuration $|0110\rangle$ is not sampled, one can construct neither eigenfunction of total spin, leading to spin contamination or redundancy (i.e. the configuration $|1001\rangle$ is involved in a CI calculation, but has coefficient 00 in the CI vector). In this work, to facilitate conservation of total spin, we relied on the following procedure: instead of collecting directly $d$ i.i.d. samples from $\mathcal{X}$ to make the batch $\mathcal{S}^{(k)}$, we collect $\sqrt{d}/2$ samples and identify all unique configurations ${\bf{x}}^{\textrm{u}}$ (u for unique) of length $M/2$ obtained from ${\bf{x}}_{\uparrow(i)}$ and ${\bf{x}}_{\downarrow(i)}$ for $21\leq i\leq\sqrt{d}/$ 2, forming the set $\mathcal{U}^{(k)}=\left\{{\bf{x}}^{\textrm{u}}\right\}$. The size of the set is upper bounded by $\left|\mathcal{U}^{(k)}\right|\leq\sqrt{d}$. From $\mathcal{U}^{(k)}$ we obtain the batch set $\mathcal{S}^{(k)}$ as:
$\mathcal{S}^{(k)}=\left\{{\bf{x}}\;\bigg{|}\;{\bf{x}}={\bf{x}}^{\textrm{u}}_{(i)}\oplus{\bf{x}}^{\textrm{u}}_{(j)}\textrm{ for all }{\bf{x}}^{\textrm{u}}_{(i)},{\bf{x}}^{\textrm{u}}_{(j)}\in\mathcal{U}^{(k)}\right\}.$ (S7)
@@
$\left(H+\lambda\big{[}S^{2}-s(s+1)\big{]}^{2}\right)|\psi\rangle=E|\psi\rangle,$ (S8)
@@
$\hat{H}_{\mathcal{S}^{(k)}}=\hat{P}_{\mathcal{S}^{(k)}}\hat{H}\hat{P}_{\mathcal{S}^{(k)}}\textrm{, with }\hat{P}_{\mathcal{S}^{(k)}}=\sum_{{\bf{x}}\in{\mathcal{S}^{(k)}}}|{\bf{x}}\rangle\langle{\bf{x}}|.$ (S9)
@@
$\left|\psi^{(k)}\right\rangle=\sum_{{\bf{x}}\in\mathcal{S}^{(k)}}c^{(k)}_{{\bf{x}}}|{\bf{x}}\rangle.$ (S10)
@@
$E^{(k)}=\left\langle\psi^{(k)}\right|\hat{H}_{\mathcal{S}^{(k)}}\left|\psi^{(k)}\right\rangle.$ (S11)
II.1.2 Self-consistent configuration recovery
@@
$\tilde{\mathcal{X}}=\left\{{\bf{x}}\;|\;{\bf{x}}\sim\tilde{P}_{\Psi}\right\}.$ (S12)
@@
$\begin{split}\sum_{p=1}^{N_{\mathrm{MO}}}\sum_{\sigma}\hat{n}_{p\sigma}|\Psi\rangle&=N|\Psi\rangle,\quad\sum_{p=1}^{N_{\mathrm{MO}}}\hat{n}_{p\uparrow}|\Psi\rangle=N_{\uparrow}|\Psi\rangle,\quad\sum_{p=1}^{N_{\mathrm{MO}}}\hat{n}_{p\downarrow}|\Psi\rangle=N_{\downarrow}|\Psi\rangle.\\ \end{split}$ (S13)
From the measurements on the quantum processors, we observe that there are a number of configurations in $\tilde{\mathcal{X}}$ whose $N_{{\bf{x}}\uparrow}$ and $N_{{\bf{x}}\downarrow}$ do not match the $N_{\uparrow}$ and $N_{\downarrow}$ of the ground state. In Table 2 we report typical values of the fraction of sampled configurations with the wrong particle number. Since the circuits we use to produce $|\Psi\rangle$ are particle-number preserving, we are certain that configurations with wrong particle numbers have been corrupted by noise. It is this subset of configurations that the configuration recovery scheme is applied to. The configurations in $\tilde{\mathcal{X}}$ whose $N_{{\bf{x}}\uparrow}=N_{\uparrow}$ and $N_{{\bf{x}}\downarrow}=N_{\downarrow}$ are not subject to the configuration recovery subroutine.
We then probabilistically flip bits and restore the correct particle number using information obtained from observables of the system. We use the spin-orbital occupancy averaged over all collected batches of subsamples ${\bf{n}}$, whose components are given by:
$n_{p\sigma}=\frac{1}{K}\sum_{1\leq k\leq K}\left\langle\psi^{(k)}\right|\hat{n}_{p\sigma}\left|\psi^{(k)}\right\rangle,$ (S14)
where $|\psi^{(k)}\rangle$ is defined in Eq. (S10). Consider a configuration ${\bf{x}}$ with $N_{{\bf{x}}\uparrow}$ spin-up electrons, where $N_{{\bf{x}}\uparrow}>N_{\uparrow}$. From the set of occupied spin-orbitals in the first half of ${\bf{x}}$, $|N_{{\bf{x}}\uparrow}-N_{\uparrow}|$ bits are sampled to be flipped. If instead $N_{{\bf{x}}\uparrow}<N_{\uparrow}$, the bits to be flipped are instead sampled from the unoccupied spin-orbitals. The same procedure applies to the spin-down orbitals.
@@
$w(y)=\begin{cases}\delta\frac{y}{h}\quad\text{if $y\leq h$},\\ \delta+(1-\delta)\frac{y-h}{1-h}\quad\text{if $y>h$}.\end{cases}$ (S15)
The parameter $h\in(0,1)$ defines the location of the “corner” of the ReLU function, while the parameter $\delta\in[0,c)$ defines the value of the ReLU function at the corner. $w$ becomes a true ReLU function when $\delta=0$, and for values of $\delta>0$ the ReLU is modified so that it is not identically zero except at $y=0$: a plot is shown in Fig. S1. In the specific cases in this work, we chose the values $\delta=0.01$ and $h=N/M$ (the filling factor) in all experiments.
Figure S1: Probability weight associated with flipping a bit with value 0 or 1 as a function of $n_{p_{\sigma}}$, following the definition in Eq. (S15). We set $\delta=0.01$ and $h=0.8$. The dashed back lines show the same probability weight associated to the naive approach of flipping according to $|x_{p\sigma}-n_{p\sigma}|$.
@@
      Obtain batches of samples $\left(\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(K)}\right)$ from $\mathcal{X}_{N}$ as described in Sec. II.1.1.
@@
      Run the eigenstate solver on the batches and obtain approximate eigenstates $\left|\psi^{(1)}\right\rangle,\ldots,\left|\psi^{(K)}\right\rangle$ (Eqs.S9 and S10).
@@
      From the approximate eigenstates construct the first guess for ${\bf{n}}$, according to Eq. (S14).
@@
      From $\mathcal{X}_{\textrm{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$, batches of samples $\left(\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(K)}\right)$ are obtained as described in Sec. II.1.1.
@@
      Run the eigenstate solver on the batches and obtain approximate eigenstates $\left|\psi^{(1)}\right\rangle,\ldots,\left|\psi^{(K)}\right\rangle$ (Eqs.S9 and S10).
@@
      From the approximate eigenstates construct refined guess for ${\bf{n}}$, according to Eq. (S14).
@@
We direct the reader to Fig. 2 in the main text for a numerical emulation of the effect of the configuration recovery on the accuracy of the HPC quantum estimator in the presence of noise.
An analytical lower bound to the probability of recovering families of configurations.–
  • $\varepsilon$; if $p$ is in the 1-sector of ${\bf{x}}_{r}$ and $x_{p}=1$.
  • $1-\varepsilon$; if $p$ is in the 1-sector of ${\bf{x}}_{r}$ and $x_{p}=0$.
  • $\varepsilon$; if $p$ is in the 0-sector of ${\bf{x}}_{r}$ and $x_{p}=0$.
  • $1-\varepsilon$; if $p$ is in the 0-sector of ${\bf{x}}_{r}$ and $x_{p}=1$.
We use these average values for our analytical derivations.
$\widetilde{\rho}=p^{L}\rho+\sigma\leavevmode\nobreak\ ,$ (S16)

### v3

@@ Results
In this study, we perform an approximate restoration of the total spin symmetry labelled by the $\hat{S}^{2}$ quantum number. In particular, we are interested on wavefunctions $|\psi^{(k)}\rangle$ that are as close as possible to singlet states, i.e., eigenfunctions of total spin with eigenvalue 0, $\hat{S}^{2}|\psi^{(k)}\rangle=0$. Conservation of symmetries is notoriously important because molecular eigenstates are joint eigenfunctions of $\hat{H}$ and its symmetries, including $\hat{S}_{z}$, $\hat{N}$, molecular point-group symmetries (if any), and $\hat{S}^{2}$. Conservation of $\hat{S}_{z}$ and $\hat{N}$ (and molecular point-group symmetries that are isomorphic to $\mathbb{Z}_{2}^{\times n}$) can be achieved in a relatively easy way, since the eigenfunctions of these operators are Slater determinants (SD) $|{\bf{x}}\rangle$, e.g.
$\hat{N}|{\bf{x}}\rangle=(N_{\alpha{\bf{x}}}+N_{\beta{\bf{x}}})|{\bf{x}}\rangle\;,\;\hat{S}_{z}|{\bf{x}}\rangle=(N_{\alpha{\bf{x}}}-N_{\beta{\bf{x}}})|{\bf{x}}\rangle\;.$ (10)
@@
Therefore, in this work, we ensure that the configurations used to span the subspace all have the desired eigenvalue of $\hat{S}_{z}$ and $\hat{N}$ as part of the configuration recovery procedure in the presence of noise. Conservation of $\hat{S}^{2}$ is more difficult to achieve in CI methods since the eigenfunctions of $\hat{S}^{2}$ are not SDs. The proper way of ensuring conservation of $\hat{S}^{2}$ is expanding $|\psi^{(k)}\rangle$ on a set of configuration state functions (CSFs), i.e. spin-symmetry-adapted linear combinations of SDs. The use of CSFs was common in early CI codes since, in addition to the obvious benefits of reducing the memory footprint of the CI vector, automatic conservation of total spin enhanced the stability of the CI iterations. Modern CI codes tend to employ SDs as opposed to CSFs, because the formation of the $\sigma$ vector (the most memory- and rate-limiting step of CI algorithms) is considerably more efficient and easily parallelized in SD-based algorithms, and in part because system memory and disk are more plentiful than in previous machines. However, in SD-based CI codes, the conservation of total spin is no longer guaranteed. In the case of our HPC quantum estimator, sampling from a quantum computer may return sets of SDs that do not allow constructing eigenfunctions of total spin. For example, in a (2e,2o) system, one may sample the configuration $|1001\rangle$ (having a single spin-down excitation over the RHF state $|0101\rangle$) which is a linear combination of the open-shell singlet and triplet states, respectively $\left(|1001\rangle\pm|0110\rangle\right)/\sqrt{2}$. If the configuration $|0110\rangle$ is not sampled, one can construct neither eigenfunction of total spin, leading to spin contamination or redundancy (i.e. the configuration $|1001\rangle$ is involved in a CI calculation, but has coefficient 00 in the CI vector). In this work, to facilitate conservation of total spin, we relied on the following procedure: instead of collecting directly $d$ independent and identically distributed (i.i.d.) samples from $\mathcal{X}$ to make the batch $\mathcal{S}^{(k)}$, we collect $\sqrt{d}/2$ samples and identify all unique configurations ${\bf{x}}^{\textrm{u}}$ (u for unique) of length $M/2$ obtained from ${\bf{x}}_{\uparrow(i)}$ and ${\bf{x}}_{\downarrow(i)}$ for $21\leq i\leq\sqrt{d}/$ 2, forming the set $\mathcal{U}^{(k)}=\left\{{\bf{x}}^{\textrm{u}}\right\}$. The size of the set is upper bounded by $\left|\mathcal{U}^{(k)}\right|\leq\sqrt{d}$. From $\mathcal{U}^{(k)}$ we obtain the batch set $\mathcal{S}^{(k)}$ as:
$\mathcal{S}^{(k)}=\left\{{\bf{x}}\;\bigg{|}\;{\bf{x}}={\bf{x}}^{\textrm{u}}_{(i)}\oplus{\bf{x}}^{\textrm{u}}_{(j)}\textrm{ for all }{\bf{x}}^{\textrm{u}}_{(i)},{\bf{x}}^{\textrm{u}}_{(j)}\in\mathcal{U}^{(k)}\right\}.$ (11)
@@
$\left(H+\lambda\big{[}S^{2}-s(s+1)\big{]}^{2}\right)|\psi\rangle=E|\psi\rangle,$ (12)
@@
$\left|\psi^{(k)}\right\rangle=\sum_{{\bf{x}}\in\mathcal{S}^{(k)}}c^{(k)}_{{\bf{x}}}|{\bf{x}}\rangle.$ (14)
@@
$E^{(k)}=\left\langle\psi^{(k)}\right|\hat{H}_{\mathcal{S}^{(k)}}\left|\psi^{(k)}\right\rangle.$ (15)
Self-consistent configuration recovery
@@
$\tilde{\mathcal{X}}=\left\{{\bf{x}}\;|\;{\bf{x}}\sim\tilde{P}_{\Psi}\right\}.$ (16)
@@
$\begin{split}\sum_{p=1}^{N_{\mathrm{MO}}}\sum_{\sigma}\hat{n}_{p\sigma}|\Psi\rangle&=N|\Psi\rangle,\quad\sum_{p=1}^{N_{\mathrm{MO}}}\hat{n}_{p\uparrow}|\Psi\rangle=N_{\uparrow}|\Psi\rangle,\quad\sum_{p=1}^{N_{\mathrm{MO}}}\hat{n}_{p\downarrow}|\Psi\rangle=N_{\downarrow}|\Psi\rangle.\\ \end{split}$ (17)
From the measurements on the quantum processors, we observe that there are a number of configurations in $\tilde{\mathcal{X}}$ whose $N_{{\bf{x}}\uparrow}$ and $N_{{\bf{x}}\downarrow}$ do not match the $N_{\uparrow}$ and $N_{\downarrow}$ of the ground state. In Table S1 we report typical values of the fraction of sampled configurations with the wrong particle number. Since the circuits we use to produce $|\Psi\rangle$ are particle-number preserving, we are certain that configurations with wrong particle numbers have been corrupted by noise. It is this subset of configurations that the configuration recovery scheme is applied to. The configurations in $\tilde{\mathcal{X}}$ whose $N_{{\bf{x}}\uparrow}=N_{\uparrow}$ and $N_{{\bf{x}}\downarrow}=N_{\downarrow}$ are not subject to the configuration recovery subroutine.
We then probabilistically flip bits and restore the correct particle number using information obtained from observables of the system. We use the spin-orbital occupancy averaged over all collected batches of subsamples ${\bf{n}}$, whose components are defined in Eq. 4.
Consider a configuration ${\bf{x}}$ with $N_{{\bf{x}}\uparrow}$ spin-up electrons, where $N_{{\bf{x}}\uparrow}>N_{\uparrow}$. From the set of occupied spin-orbitals in the first half of ${\bf{x}}$, $|N_{{\bf{x}}\uparrow}-N_{\uparrow}|$ bits are sampled to be flipped. If instead $N_{{\bf{x}}\uparrow}<N_{\uparrow}$, the bits to be flipped are instead sampled from the unoccupied spin-orbitals. The same procedure applies to the spin-down orbitals.
@@
$w(y)=\begin{cases}\delta\frac{y}{h}\quad\text{if $y\leq h$},\\ \delta+(1-\delta)\frac{y-h}{1-h}\quad\text{if $y>h$}.\end{cases}$ (18)
The parameter $h\in(0,1)$ defines the location of the “corner” of the ReLU function, while the parameter $\delta\in[0,c)$ defines the value of the ReLU function at the corner. $w$ becomes a true ReLU function when $\delta=0$, and for values of $\delta>0$ the ReLU is modified so that it is not identically zero except at $y=0$ In the specific cases in this work, we chose the values $\delta=0.01$ and $h=N/M$ (the filling factor) in all experiments.
@@
      Obtain batches of samples $\left(\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(K)}\right)$ from $\mathcal{X}_{N}$ as described in Sec. Eigenstate solver.
@@
      Run the eigenstate solver on the batches and obtain approximate eigenstates $\left|\psi^{(1)}\right\rangle,\ldots,\left|\psi^{(K)}\right\rangle$ (Eqs.13 and 14).
@@
      From the approximate eigenstates construct the first guess for ${\bf{n}}$, according to Eq. (4).
@@
      From $\mathcal{X}_{\textrm{R}}=\mathcal{X}_{N}\cup\mathcal{X}_{\rightarrow N}$, batches of samples $\left(\mathcal{S}^{(1)},\ldots,\mathcal{S}^{(K)}\right)$ are obtained as described in Sec. Eigenstate solver.
@@
      Run the eigenstate solver on the batches and obtain approximate eigenstates $\left|\psi^{(1)}\right\rangle,\ldots,\left|\psi^{(K)}\right\rangle$ (Eqs.13 and 14).
@@
      From the approximate eigenstates construct refined guess for ${\bf{n}}$, according to Eq. (4).
@@
We direct the reader to Fig. 2 in the main text for a numerical emulation of the effect of the configuration recovery on the accuracy of SQD in the presence of noise.
Energy-variance extrapolation
It is expected that the accuracy of SQD (with or without configuration recovery) will increase as the number of configurations used for the subspace expansion $d$ is increased. However, the convergence of the ground-state properties with $d$ is not expected to follow any specific functional relation. Therefore, attempting to analyze the convergence as a function of $d$ is not well motivated. Instead, the different eigenstate approximations obtained for different values of $d$ and different batches of samples $\mathcal{S}^{(k)}$ are used for an energy-variance extrapolation  (?, ?, ?, ?, ?, ?). Consider the approximate eigenstate $\left|\psi\right\rangle$, whose energy is given by $E=\langle\psi|\hat{H}|\psi\rangle$, and consider the exact eigenstate energy $E_{\textrm{T}}$. The difference between the approximate energy and $E_{\textrm{T}}$,
$\delta E=\langle\hat{H}\rangle-E_{\textrm{T}},$ (19)
vanishes linearly with the Hamiltonian variance divided by the square of the variational energy  (?)
$\frac{\Delta H}{E^{2}}=\frac{\langle\psi|\hat{H}^{2}|\psi\rangle-\langle\psi|\hat{H}|\psi\rangle^{2}}{E^{2}}.$ (20)
In the main text and in this supplementary material we apply the energy-variance analysis to two different sets of energy-variance pairs. The first one is the energy-variance pairs obtained by the HCI procedure where different energies and variances are obtained by changing the cutoff parameter that indirectly controls the number of determinants in the subspace projection and diagonalization, resulting in different levels of accuracy. The second set of energy-variance pairs are those obtained from SQD for different numbers of configurations $d$ as well as for different batches of sampled configurations $\mathcal{S}^{(k)}$.
Experimental details
In this section we describe the quantum circuits that are employed to produce the configurations to which we apply the eigenstate solver. Additionally, we provide details on the setting of circuit parameters from an efficient classical CCSD calculation, and the mapping of the circuits to quantum processors with heavy-hex connectivity.
Quantum circuits
In this work, we used the local-unitary cluster Jastrow (LUCJ) ansatz  (?) to sample randomly distributed electronic configurations. LUCJ derives from the unitary cluster Jastrow (UCJ) ansatz, which has the form  (?) of a product of $L$ layers, as defined in Eq. 5. We recall that:
$\hat{K}_{\mu}=\sum_{pq,\sigma}K_{pq}^{\mu}\,\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{q\sigma}\;,\;\hat{J}_{\mu}=\sum_{pr,\sigma\tau}J_{p\sigma,q\tau}^{\mu}\,\hat{n}_{p\sigma}\hat{n}_{q\tau}\;.$ (21)
In Eq. (21), $p,q=0\dots N_{\mathrm{MO}}-1$ label molecular spatial orbitals and $\sigma,\tau$ label spin polarizations ($\alpha,\beta$ for spin-up and spin-down electrons respectively). $K_{pq}^{\mu}$/$J_{pq,\sigma\tau}^{\mu}$ has complex/real matrix elements and is anti-Hermitian/symmetric. The UCJ ansatz can be derived from a twice-factorized low-rank decomposition of the qUCCD ansatz  (?, ?), and the $L$-product form is such that the exact full configuration interaction wavefunction can be obtained via Eq. (5)  (?, ?).
The local UCJ or LUCJ  (?) introduces a “local” approximation of the UCJ ansatz, which makes the following modifications for opposite-spin and same-spin number-number terms:
$\begin{split}\sum_{pq}J_{p\alpha,q\beta}\hat{n}_{p\alpha}\hat{n}_{q\beta}&\rightarrow\sum_{p\in S}J_{p\alpha,p\beta}\hat{n}_{p\alpha}\hat{n}_{p\beta}\\ \sum_{pq}J_{p\sigma,q\sigma}\hat{n}_{p\sigma}\hat{n}_{q\sigma}&\rightarrow\sum_{pq\in S^{\prime}}J_{p\sigma,q\sigma}\hat{n}_{p\sigma}\hat{n}_{q\sigma},\end{split}$ (22)
where $\sigma=\alpha,\beta$ and the sets $S,S^{\prime}$ are such that the quantum circuit implementing $e^{i\hat{J}_{\mu}}$ has depth $O(1)$ and only comprises $O(|S|+|S^{\prime}|)$ number-number “nn gates,” i.e. two-qubit unitaries of the form $U_{nn}(\varphi)=e^{-i\frac{\varphi}{4}(Z_{p}+Z_{q}-Z_{p}Z_{q})}$, acting on adjacent qubits $p,q$ in the topology of a certain processor. For example, on a heavy-hex processor, $S=\{4k\;,\;k=0\dots(N_{\mathrm{MO}}-1)/4\}$ and $S^{\prime}=\{(p,p+1)\;,\;p=0\dots N-2\}$  (?). The circuits $e^{\pm\hat{K}_{\mu}}$, on the other hand, can be implemented by a Bogolyubov circuit acting on $N_{\mathrm{MO}}$ qubits and comprising $O(N_{\alpha}(N_{\mathrm{MO}}-N_{\alpha}))$ gates and depth $O(N_{\mathrm{MO}})$  (?, ?, ?). Through the local approximation, the LUCJ Ansatz balances hardware friendliness and accuracy, the latter ultimately deriving from its connection to coupled cluster theory and adiabatic state preparation  (?).
Initialization of LUCJ parameters